# Direction-Dependent Conduction Polarity in Altermagnetic CrSb

### v1

@@ 3 Conclusion
In summary, we have studied the anisotropic transport properties of altermagnetic CrSb. Both Hall and Seebeck coefficients indicate that CrSb exhibits DDCP, with in-plan conduction dominated by electrons and cross-plane conduction dominated by holes. Our DFT calculations based on the Boltzmann transport equation and the relaxation-time approximation qualitatively support our experimental results. We identify two hole pockets and one electron pocket dominating the cross-plane and in-plane conduction, respectively. Furthermore, our calculations predict that the DDCP in CrSb is fragile and can be suppressed with a small amount of doping. This prediction is validated by the loss of DDCP in minimally hole doped Cr0.98V0.02Sb. Composed of earth-abundant and non-toxic elements, CrSb emerges as a promising candidate for practical applications leveraging the DDCP behavior. Furthermore, the inherent altermagnetic nature of CrSb, characterized by large spin-splitting, opens up exciting possibilities for future spintronic applications that can exploit both DDCP and spin-splitting phenomena.
@@ 4 Method
The thermopower measurements were carried out on a bulk single crystal in both directions. The two probe electrodes were made on top of the crystal sample by fixing Cu-wires using silver paste. These measurements were done in a liquid nitrogen cryostat where the sample was kept under vacuum. The temperature difference between two ends of the sample was generated by sending currents from a source meter (Keithley 2450) to the heater attached just beside the sample. The voltage and temperature difference between two ends of the sample were measured by a precision nano voltmeter (Keithley 2182A) and temperature controller (Lakeshore 340). The bath temperature is controlled using a temperature controller (Lakeshore 340) with temperature fluctuations less than ±$\pm 20\,\mathrm{mK}$.
DFT Calculations: The electronic structure was computed within the Vienna Ab initio Simulation Package (VASP), employing a plane-wave basis set and projected augmented wave potentials.[28, 29, 30, 31] The generalized gradient approximation (GGA) [32] for the exchange-correlation functional was used. While the lattice parameters were fixed at the experimental values [33] for the NiAs structure of CrSb in the P63/mmc spacegroup, all internal atomic positions were optimized through a total energy minimization. A cutoff energy of 500 eV was used to define the maximum kinetic energy for the plane waves in the basis. Brillouin zone integrations were performed with a Monkhorst-Pack k-grid of density $10\times 10\times 10$ points.[34] The Seebeck and Hall coefficients were calculated using the semiclassical Boltzmann transport equation. These calculations employed the relaxation time and rigid band approximations, with the DFT band dispersions interpolated on a dense 35×35×35 k-mesh grid as implemented in Boltztrap2.[35] Additionally, the inverse effective mass (IEM) tensor was determined using B-spline interpolation of the DFT band dispersions.
@@ Acknowledgements
This work utilized the instumentation facilities provided by the Technical Research Centre (TRC) at S. N. Bose National Centre for Basic Sciences, under Department of Science and Technology (DST), Government of India. NK acknowledges the Science and Engineering Research Board (SERB), India, for financial support through Grant Sanction No. CRG/2021/002747 and Max Planck Society for funding under Max Planck-India partner group project. PM acknowledges support from SERB through the project SERB-POWER (SPF/2021/000066). MM acknowledges CSIR-Human Resource Development Group (HRDG) (03/1511/23/EMR-II). BR acknowledges the financial support from DST.

### v2

@@ 3 Conclusion
In summary, we have studied the anisotropic transport properties of altermagnetic CrSb. Both Hall and Seebeck coefficients indicate that CrSb exhibits DDCP, with in-plane conduction dominated by electrons and cross-plane conduction dominated by holes. Our DFT calculations based on the Boltzmann transport equation and the relaxation-time approximation qualitatively support our experimental results. We show that the DDCP in CrSb arises from the multicarrier mechanism and identify two hole pockets and one electron pocket dominating the cross-plane and in-plane conduction, respectively. The emergence of DDCP in CrSb is a consequence of its specific Fermi surface geometry, which is shaped by altermagnetic spin splitting. However, it is important to emphasize that altermagnetic spin splitting does not inherently lead to DDCP in all materials. In CrSb, the particular Fermi surface geometry happens to facilitate DDCP, but this behavior is not a universal feature of altermagnets. Furthermore, our calculations predict that the DDCP in CrSb is fragile and can be suppressed with a small amount of doping. This prediction is validated by the loss of DDCP in minimally hole doped Cr0.98V0.02Sb. Composed of earth-abundant and non-toxic elements, CrSb emerges as a promising candidate for practical applications leveraging the DDCP behavior. Furthermore, the inherent altermagnetic nature of CrSb, characterized by large spin-splitting, opens up exciting possibilities for future spintronic applications that can exploit both DDCP and spin-splitting phenomena.
@@ 4 Method
The thermopower measurements were carried out on a bulk single crystal in both directions. The two probe electrodes were made on top of the crystal sample by fixing Cu-wires using silver paste. These measurements were performed in a liquid nitrogen cryostat where the sample was kept under vacuum. The temperature difference between two ends of the sample was generated by sending currents from a source meter (Keithley 2450) to the heater attached just beside the sample. The voltage and temperature difference between two ends of the sample were measured by a precision nano voltmeter (Keithley 2182A) and temperature controller (Lakeshore 340). The bath temperature is controlled using a temperature controller (Lakeshore 340) with temperature fluctuations less than ±$\pm 20\,\mathrm{mK}$.
DFT Calculations: The electronic structure was computed within the Vienna Ab initio Simulation Package (VASP), employing a plane-wave basis set and projected augmented wave potentials.[32, 33, 34, 35] The generalized gradient approximation (GGA) [36] for the exchange-correlation functional was used. While the lattice parameters were fixed at the experimental values [37] for the NiAs structure of CrSb in the $P6_{3}/mmc$ space group, all internal atomic positions were optimized through a total energy minimization. A cutoff energy of 500 eV was used to define the maximum kinetic energy for the plane waves in the basis. Brillouin zone integrations were performed with a Monkhorst-Pack k-grid of density $10\times 10\times 10$ points.[38] The Seebeck and Hall coefficients were calculated using the semiclassical Boltzmann transport equation. These calculations employed the relaxation time and rigid band approximations, with the DFT band dispersions interpolated on a dense 35×35×35 k-mesh grid as implemented in Boltztrap2.[30] Additionally, the inverse effective mass (IEM) tensor was determined using B-spline interpolation of the DFT band dispersions.
@@ Acknowledgements
This work utilized the instrumentation facilities provided by the Technical Research Centre (TRC) at S. N. Bose National Centre for Basic Sciences, under Department of Science and Technology (DST), Government of India. NK acknowledges the Science and Engineering Research Board (SERB), India, for financial support through Grant Sanction No. CRG/2021/002747 and Max Planck Society for funding under Max Planck-India partner group project. PM acknowledges support from SERB through the project SERB-POWER (SPF/2021/000066). MM acknowledges CSIR-Human Resource Development Group (HRDG) (03/1511/23/EMR-II). BR acknowledges the financial support from DST. KP acknowledges Param Rudra Facility under the National Supercomputing Mission, Government of India at S. N. Bose National Centre for Basic Sciences.
@@
Supporting Information
Direction-Dependent Conduction Polarity in Altermagnetic CrSb
Banik Rai1,3, Krishnendu Patra1,3, Satyabrata Bera2, Sk Kalimuddin2, Kakan Deb1,
Mintu Mondal2, Priya Mahadevan1, Nitesh Kumar1
1Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India
2School of Physical Sciences, Indian Association for the Cultivation of Science, Jadavpur, Kolkata-700032, India
3These authors contributed equally.
S1      EDXS and XRD
Figure S1 shows the typical EDXS spectrum of the elemental composition of CrSb and Cr0.98V0.02Sb. The EDXS results, obtained from several spots on single crystals of CrSb and Cr0.98V0.02Sb, are tabulated in Table S1. The elemental ratio for CrSb is close to 1:1. Upon doping with V, the nominal EDXS composition is nearly Cr0.98V0.02Sb.
Figure S1: EDXS spectrum of CrSb (blue) and Cr0.98V0.02Sb (red).
Table S1: Chemical composition and derived chemical formulas.
Compound %Cr %V %Sb %$\langle\%Cr\rangle$$\langle\%V\rangle$$\langle\%Sb\rangle$ Chemical Formula
CrSb 49.1 - 50.1 50.325 - 49.675 Cr1.0065Sb0.9935
50.3 - 49.7
51.0 - 49.0
50.9 - 49.1
Cr0.98V0.02Sb 49.5 0.8 49.7 49.26 0.84 49.86 Cr0.985(6)V0.017(5)Sb0.997(2)
49.1 1.0 49.9
49.6 0.5 49.9
48.8 1.2 49.9
49.3 0.7 49.9
Figure S2(a) shows the powder XRD pattern of CrSb fitted with the space group $P63/mmc$ using the Le-bail method. Fig. S2(b) shows the XRD pattern of a plate-like single crystal of CrSb. The observation of only $(000l)$ peaks confirms that the surface exposed to the XRD beam is ab-plane and the absence of other peaks indicates good crystallinity of the single crystal.
Figure S2: XRD study on CrSb. (a) Powder XRD pattern of CrSb fitted with space group $P6_{3}/mmc$ using the Le-bail method. (b) XRD pattern of a plate-like single crystal of CrSb.
S2      Magnetoresistance and multi-carrier fitting
Figure S3(a)-(c) shows the magnetic field-dependent transverse magnetoresistance $\bigg{[}\mathrm{MR}=\dfrac{\rho_{ii}(B)-\rho_{ii}(0)}{\rho_{ii}(0)}\times 100\>\%\bigg{]}$ of CrSb at different temperatures, measured along the three orthogonal x, y, and z directions. MR reaches its maximum value of 8% for $B||\mathrm{z}$ [Figure S3(c)] at 9 T and 2 K. For the other two directions, the MR is smaller, with the smallest value observed for $B||\mathrm{x}$ [Fig. S3(a)]. In all directions, the observed positive MR can be attributed to the orbital cyclotron motion of charge carriers under the applied magnetic field. In systems with nearly equal electron and hole carrier densities (charge compensation), the MR typically exhibits a quadratic dependence on the magnetic field.[1] Figure S3(d), showing the MR at 2 K for all three directions, highlights the deviation of the MR from the quadratic behavior, indicating the presence of uncompensated charge carriers in CrSb.
In materials with multiple bands crossing the Fermi level ($E_{\mathrm{F}}$), the mobility ($\mu$) of charge carriers often varies between the bands. As a result, the charge carriers move independently within these bands, each characterized by its own carrier density ($n_{i}$) and mobility ($\mu_{i}$).[2] This results in a non-linear Hall signal as a function of the magnetic field. The longitudinal ($\sigma_{\mathrm{xx}}$) and Hall ($\sigma_{\mathrm{xy}}$) conductivity, considering the multi-carrier effect, can be expressed as:
Figure S3: (a)-(c) Magnetoresistance of CrSb measured in various geometries at different temperatures. (d) Magnetoresistance at 2K measured along various directions and the corresponding exponential fit.
$\sigma_{\mathrm{xx}}$ $=e\sum\frac{n_{i}\mu_{i}}{1+\mu_{i}^{2}B^{2}}$ (S1)
$\sigma_{\mathrm{xy}}$ $=eB\sum\frac{n_{i}\mu_{i}^{2}}{1+\mu_{i}^{2}B^{2}}.$ (S2)
Table S2: Carrier density and mobility at various temperatures obtained by three carrier fit.
Temperature $n_{\mathrm{h}}$ ${n}_{\mathrm{e1}}$ ${n}_{\mathrm{e2}}$ $\mu_{\mathrm{h}}$ $\mu_{\mathrm{e1}}$ $\mu_{\mathrm{e2}}$
(K) $(10^{21}\mathrm{cm}^{-3})$ $(10^{20}\mathrm{cm}^{-3})$ $(10^{19}\mathrm{cm}^{-3})$ $(\mathrm{cm}^{2}\mathrm{V}^{-1}\mathrm{s}^{-1})$ $(\mathrm{cm}^{2}\mathrm{V}^{-1}\mathrm{s}^{-1})$ $(\mathrm{cm}^{2}\mathrm{V}^{-1}\mathrm{s}^{-1})$
2 1.45 6.39 1.75 312 480 2394
10 1.42 6.61 1.84 316 471 2344
20 1.39 6.78 1.87 315 458 2304
50 1.44 6.20 1.30 264 435 2196
100 1.12 7.63 0.89 204 281 1674
150 0.69 11.2 0.69 186 165 1293
Figure S4(a)-(b) shows the field-dependent $\sigma_{\mathrm{xx}}$ and $\sigma_{\mathrm{xy}}$ at 2 K fitted using equation (S1) and (S2), considering two and three carrier models. The fitting quality for two-carrier model is poor, indicating the presence of more than two independent types of charge carriers. The fitting improves significantly with the three-carrier model. This model also fits the data at higher temperatures, as shown in Figure S4(c)-(d). The values of $n$ and $\mu$ obtained from the fit are presented in Figures S4(e) and S4(f), respectively. Of the three independent charge carriers, one carrier type is hole-like (h) and the other two are electron-like (e1 and e2). The hole-like carrier has the highest carrier density ($n_{h}\approx 1\times 10^{21}\mathrm{cm}^{-3}$). The two electron-like carriers have smaller carrier density, with one being much smaller than the other. The mobilities follow the opposite trend. The hole mobility $(\mu_{h})$ is the smallest and the mobility of the second electron-like carrier ($\mu_{e2}$) is the largest, reaching a value as large as 2394 $\mathrm{cm^{2}V^{-1}s^{-1}}$ at 2 K. Table S2 shows the variation of carrier density and mobility of three carriers with temperature as obtained from the three-carrier fit.
Figure S4: Multicarrier effect in the transport of CrSb. (a)-(b) Global (simultaneous) fitting of longitudinal ($\sigma_{\mathrm{xx}}$) and Hall ($\sigma_{\mathrm{xy}}$) conductivity at 2 K using two-carrier and three-carrier model. (c)-(d) Global fitting of $\sigma_{\mathrm{xx}}$ and $\sigma_{\mathrm{xy}}$ at various temperatures using three-carrier model. Variation of (e) carrier density and (f) mobility with temperature as obtained from three carrier fit. These values are in close agreement with the values reported in Ref. [3].
Figure S5: The attribution of different carrier types, as obtained from the three-carrier fit, to different portions of the FS.
The Fermi surface (FS) of CrSb consists of three pockets, namely $\alpha$, $\beta$, and $\gamma$, corresponding to three distinct bands crossing $E_{\mathrm{F}}$ (see main text). Effective mass calculations (main text Table 1) reveal that the $\alpha$ and $\beta$ are hole pockets, while $\gamma$ is an electron pocket. This seemingly contradicts the results of the three-carrier fitting, which identifies two electron-like and one hole-like carriers. However, this apparent discrepancy can be resolved by considering the origin of the $\alpha$ and $\beta$ pockets. These pockets arise from the altermagnetic and spin-orbit splitting of the same energy band. Consequently, they exhibit strong similarities in their electronic properties. Furthermore, the effective masses of the charge carriers residing within these pockets are nearly identical, suggesting that their mobilities are also highly comparable. Due to these strong similarities, the charge carriers within the $\alpha$ and $\beta$ pockets do not behave as entirely independent entities in transport. Their contributions tend to be highly correlated, effectively merging into a single, dominant hole-like carrier contribution in the three-carrier fit. The first electron-like carrier (e1) in the three-carrier fit can be attributed to the electron pocket $\gamma$. The second electron-like carrier (e2) most likely correspond to the small hemispherical sub-pockets present within the $\alpha$ and $\beta$ pockets, as these sub-pockets are electron-like in nature (see main text). Since, these sub-pockets are well separated from the larger portion of the $\alpha$ (or $\beta$) pocket and exhibit distinct geometry, the charge carriers residing on them can contribute independently to the transport, even though they arise from the same energy band. The value of ${n_{e2}}$ is about 100 times smaller than ${n_{h}}$, as evident from Table S2. A similar proportion is visibly evident in the size of the larger portion of the $\alpha$ (or $\beta$) pocket and its sub-pocket. Since, carrier density is proportional to the volume of the corresponding Fermi pocket, our attribution of second electron-like carrier to the hemispherical sub-pockets is well supported. The value of $n_{e1}$ is also consistent with the relative size of the $\gamma$ pocket. In Figure S4, different carrier types, as obtained from the three-carrier fit, have been attributed to specific portions of the FS.
S3     Altermagnetic splitting of energy bands
Altermagnets have been identified as a new phase of collinear magnets that share properties with both ferromagnets (e.g., non-relativistic spin splitting of energy bands) and antiferromagnets (e.g., zero net magnetization)[4]. In altermagnets, the spatial arrangement of non-magnetic atoms surrounding the magnetic ones plays a crucial role in dictating their distinct spin group symmetry, which distinguishes them from ferromagnets and conventional antiferromagnets. Unlike conventional antiferromagnets, where the opposite spin sublattices are connected by a simple lattice translation or inversion, in altermagnets, this connection is established by a real-space crystal rotation. The symmetries of an antiferromagnetic crystal, i.e. $\mathcal{T}$, ${\textit{{t}}}\mathcal{T}$ and $\mathcal{P}\mathcal{T}$, where $\mathcal{T}$, t and $\mathcal{P}$ are time reversal, lattice translation and space inversion operations, are thus broken in an altermagnet which facilitates the momentum dependent spin splitting of energy bands even without spin-orbit coupling akin to ferromagnets. For CrSb, the sublattice transposing symmetries are $C_{\mathrm{6z}}t_{1/2}\tau$ and $M_{\mathrm{z}}\tau$ (see main text). The first symmetry, $C_{\mathrm{6z}}t_{1/2}\tau$, results in three spin-degenerate nodal planes in the Brillouin zone (BZ) as shown by the blue vertical planes in Figure S6(a). The second symmetry, $M_{\mathrm{z}}\tau$, result in one spin-degenerate nodal plane in the BZ, as shown by a orange horizontal plane in Figure S6(a). Spin-degeneracy is lifted when moving away from these four nodal planes. Figure S6(b) shows the band structure of CrSb calculated along the path $\mathrm{L_{1}}-\Gamma-\mathrm{L_{2}}$ of the BZ in the absence of spin-orbit coupling (SOC). As this path lies outside the four nodal-planes, the bands experience significant splitting.
Figure S6: (a) Brillouin zone (BZ) of CrSb showing four spin-degenerate nodal planes. (b) Band structure of CrSb calculated in the absence of spin-orbit coupling (SOC) along the paths as shown in Figure S6(a). A significant altermagnetic splitting can be observed. The dashed horizontal line represents the experimental fermi energy ($E^{\prime}_{\mathrm{F}}$) (see main text).
S4      Magnetic component-projected relativistic band dispersion
Figure S7: Band dispersions for CrSb plotted along various symmetry directions. Spin-orbit interactions have been included in the calculation of the band structure. The spin projections for positive (a) $\langle\mathrm{m}_{x}\rangle$ (c) $\langle\mathrm{m}_{y}\rangle$ and (e) $\langle\mathrm{m}_{z}\rangle$ as well as negative (b) $\langle\mathrm{m}_{x}\rangle$ (d) $\langle\mathrm{m}_{y}\rangle$ and (f) $\langle\mathrm{m}_{z}\rangle$ have been shown. The color bar on the right of each plot gives the magnitude.
S5      Band dispersion at finite $k_{z}$ plane
It is evident from the FS that there is no $\gamma$ pocket at the $k_{z}=0$ plane, as there is no intersection between band #3 and the shifted Fermi energy ($E^{\prime}_{\mathrm{F}}$) (indicated by dashed horizontal line) in that plane. However, at a finite $k_{z}$, band #3 intersects $E^{\prime}_{\mathrm{F}}$ which gives rise to the $\gamma$ pocket in the FS. In Figure S7, relativistic band dispersion is shown for the $k_{z}=0.2\pi/c$ plane, where the band #3 intersects $E^{\prime}_{\mathrm{F}}$.
Figure S8: Band dispersion (with SOC turned on) along high symmetry lines for a particular fixed plane $k_{z}=0.2\pi/c$.
S6      Number of electrons per unit cell (u.c.) as a function of energy.
We have calculated the number of electrons at a particular energy using DFT calculations. From this, we determined the change in the number of electrons at the shifted Fermi energy ($E^{\prime}_{\mathrm{F}}$) as shown in Figure S9.
Figure S9: Number of electrons per unit cell as a function of energy.
S7      Seebeck coefficient of Cr0.98V0.02Sb
To prove that a small amount of hole doping in CrSb destroy the DDCP, we have measured the in-plane Seebeck coefficient ($S_{\mathrm{xx}}$) of Cr0.98V0.02Sb. CrSb is n-type in the ab-plane and thus has negative $S_{\mathrm{xx}}$. In Figure S10, we show that $S_{\mathrm{xx}}$ changes sign and becomes positive for Cr0.98V0.02Sb, thus destroying the DDCP.
Figure S10: In-plane seebeck coefficient ($S_{\mathrm{xx}}$) for CrSb and Cr0.98V0.02Sb, showing that the sign of $S_{\mathrm{xx}}$ has changed from negative to positive in Cr0.98V0.02Sb because of hole (vanadium) doping.